# Gated Rotary-Enhanced Linear Attention for Long-term Sequential Recommendation

### v1

@@ 3. METHODOLOGY
Figure 2. The overview of our proposed architecture. The Overall Architecture (left) processes an input sequence through an Embedding Layer, which is then fed into a stack of $L$ identical encoder layers. Each encoder layer consists of Layer Normalization, a GRELA Block, another Layer Normalization, and an MLP, finally leading to a Prediction Layer. The central components are the GRELA Block (middle-bottom), which utilizes RoPE (Rotary Position Encoding, middle-top), and the RELA module (right). Within the RELA module, query ($Q$) and key ($K$) are processed by ELU activation and the RoPE. Attention scores are computed via matrix multiplication of the RoPE-enhanced $Q$ and $K$ (after $K$ is scaled), followed by another scaling and a concatenation step. A Causal Conv1D and SiLU activation processes the value ($V$) projection to capture local context.
@@
The presentation is structured as follows: initially, we introduce the embedding layer; secondly, we elaborate on the core components of the GRELA Block; subsequently, we present the details of the Multi-Layer Perceptron, and finally, we describe the prediction layer and training strategies. The key notations of this section are listed in Table 2.
3.1. Embedding Layer
@@
Following (Kang and McAuley, 2018), we utilize an embedding layer to map the training sequence $\mathcal{S}_{u}$ into a high-dimensional space. The embedding layer uses a learnable embedding matrix $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $\mathcal{V}$ denotes the item set and $D$ represents the embedding dimension. Given an input interaction sequence $\mathcal{S}_{u}$ of length $N$, it can be projected into the initial embedding matrix $\boldsymbol{H}_{o}\in\mathbb{R}^{N\times D}$ as follows:
(9) $\boldsymbol{H}_{o}=\boldsymbol{E}\mathcal{S}_{u}.$
@@
Then we apply the Dropout (Srivastava et al., 2014) and layer normalization (Lei Ba et al., 2016) operations to obtain the final embedding $\boldsymbol{H}\in\mathbb{R}^{N\times D}$:
(10) $\boldsymbol{H}=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\boldsymbol{H}_{o}\right)\right).$
@@
Attention-based sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020; Liu et al., 2023) typically inject a learnable position encoding into the final embedding, while our method captures position awareness with rotary position encoding by multiplying the rotation matrix with outputs of the kernel function in linear attention. This design eliminates the need to incorporate positional information into the embedding layers.
3.2. GRELA Block
@@
3.2.1. Rotary-Enhanced Linear Attention
@@
Overall Design. Previous work (Liu et al., 2023) provides $\mathcal{O}(N)$ computational complexity of attention mechanism to alleviate the issue of quadratic complexity in dot-product attention for sequential recommendation. However, it ignores the high cost of learnable position encoding in long-sequence modeling. Meanwhile, some works (Tian et al., 2024; Lopez-Avila et al., 2024) notice these issues and propose improved RoPE to address them, but unfortunately, they fail to address the problem of dot-product attention with $\mathcal{O}(N^{2})$ complexity. To bridge this gap, we propose a novel Rotary-Enhanced Linear Attention (RELA) mechanism, aiming to simultaneously resolve the issues of quadratic complexity and ineffective position encoding in long-sequence modeling. The motivation stems from the observation that other position encoding methods operate directly on the input sequence or the attention matrix. However, linear attention does not compute the attention matrix, which means other methods can only modify the input sequence. In contrast, RoPE achieves relative position encoding by leveraging absolute position encoding, by applying transformations only to the queries $\boldsymbol{Q}$ and keys $\boldsymbol{K}$. This design makes RoPE suitable for integration with linear attention. In a general situation, following Eq.(7), we revise Eq. (3) as follows:
(11) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
@@
where $\mathcal{R}\left(\cdot\right)$ denotes injecting the RoPE into queries or keys. Following (Katharopoulos et al., 2020), we set $\phi(x)=\operatorname{ELU}(x)+1$. This kernel function ensures that the denominator remains positive to avoid the risk of division by zero, while the summation in the numerator also contains no negative terms.
@@
Inspired by techniques for stabilizing gradients in dot-product attention (Vaswani et al., 2017), we further incorporate scaling factors. Specifically, the keys $\boldsymbol{K}$ and values $\boldsymbol{V}$ are scaled by $\sqrt{N}$ (where $N$ denotes the sequence length). This emulates the scaling mechanism used in conventional dot-product attention and helps regulate the dynamic range of the resulting key-value memory. Furthermore, the normalization term in the denominator, which is critical for linear attention, is constructed using the dot product between the query $\phi(\boldsymbol{Q}_{m})$ and the mean of keys $\phi(\boldsymbol{K}_{n})$. A small constant $\epsilon$ is added to this denominator for numerical stability, preventing division by zero and ensuring well-behaved gradients. This approach enables more fine-grained control over the attention distribution. The final RELA is formulated as follows:
(12) $\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)/\sqrt{N}\right)\left(\boldsymbol{V}_{n}/\sqrt{N}\right)\right)}{\phi(\boldsymbol{Q}_{m})\left(\frac{1}{N}\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}+\epsilon\right)}.$
@@
(13) $\boldsymbol{Q}_{m}=\left[\boldsymbol{Q}_{m}^{(1)},\boldsymbol{Q}_{m}^{(2)}\right]=\left[\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right],$
@@
(14) $e^{im\theta}=\cos m\theta+i\sin m\theta.$
@@
(15) $\boldsymbol{Q}_{m}e^{im\theta}$ $=\left(\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right)(\cos(m\theta)+i\sin(m\theta))$
@@
(16) $\boldsymbol{Q}_{m}e^{im\theta}$ $=[\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta),\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)]$
@@
While the core rotational mechanism of RoPE effectively injects relative positional information, a critical design choice is the dimension upon which these rotations are applied. Conventionally, in multi-head attention architectures, RoPE is applied independently to each head’s query and key vectors, meaning the rotations operate on the head dimension $d$ (Su et al., 2024; Tian et al., 2024). This allows each head to potentially learn or leverage positional cues in its specialized subspace. Our choice deviates from this standard practice. We apply RoPE directly to the full embedding dimension $D$ of the queries and keys. The motivation is that imbuing the entire token representation with a unified positional signal at the embedding dimension $D$ level can lead to a more holistic and consistent understanding of sequence order throughout the model. Instead of each head independently re-interpreting or establishing positional context, all downstream computations operate on an input that is already position-aware across its full breadth.
3.2.2. Local Shortcut and Gated Mechanism
Attention mechanisms excel at long-range modeling, yet local behavior modeling is equally vital for long-term sequential recommendation systems, as user interactions often display strong local preferences (e.g., session-based behaviors) crucial for next-item prediction. However, linear attention variants falter here. Unlike Softmax attention’s sharp focus, linear attention’s smooth distribution tends to average features, obscuring crucial local cues (Han et al., 2023). Moreover, RoPE suffers from long-term decay (Su et al., 2024), meaning the attention score decreases as the relative distance increases. Because of linear attention’s smoothing and RoPE’s distance decay, precisely modeling crucial local patterns for long-term sequential recommendation becomes difficult. To address these issues, we introduce a Local Shortcut to mitigate the long-term decay of RELA. Specifically, we use a causal 1D convolution (Fu et al., 2023) followed by a SiLU activation (Ramachandran et al., 2017) to capture local preferences in the user interaction sequence:
(17) $\boldsymbol{S}=\operatorname{SiLU}\left(\text{CausalConv1d}\left(\boldsymbol{V}\right)\right),$
@@ 3.2.2. Local Shortcut and Gated Mechanism
(18) $\boldsymbol{O}_{f}=\boldsymbol{O}+\boldsymbol{S}.$
@@
The theoretical motivation behind such a design concerns the upper bound on the rank of the attention matrix in linear attention (Han et al., 2023). The rank of the attention matrix is bounded by the length of the sequence $N$ and the feature dimension $d$ for each head:
(19) $\operatorname{rank}(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})^{\top})$ $\leq\min\{\operatorname{rank}(\phi(\boldsymbol{Q})),\operatorname{rank}(\phi(\boldsymbol{K}))\}$
@@
(20) $\boldsymbol{O}_{f}$ $=\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\operatorname{CSC}(\boldsymbol{V})$
@@
To address the challenge of differentiating between temporary local preferences and genuine shifts in long-term user interests, we design a gated mechanism to intelligently modulate RELA’s output. We utilize a SiLU gate (Qin et al., 2023; Gu and Dao, 2024) and functioning as an output gate, learns to derive weights from the interaction sequence. These weights then adjust the influence of different signals within the user behavior sequence processed by RELA. This allows the model to discern whether a new burst of user behaviors represents a strong but potentially transient local preference or signifies a true evolution in the user’s underlying taste.
The overall structure of the GRELA Block is shown as follows:
(21) $\boldsymbol{H}_{g}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}\right),$
@@
where $\boldsymbol{H}_{g}$ is the input of the GRELA block, i.e., when $L=1$, the input $\boldsymbol{H}$ is from the embedding layer, $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{D\times D}$ are the weight matrices of the linear projection, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$ are the biases. Then a Hadamard product is performed on $\boldsymbol{H}_{1}$ and $\boldsymbol{H}_{2}$ followed by a linear projection and Droppath (Larsson et al., 2017) operation:
(22) $\boldsymbol{H}_{3}=\operatorname{DropPath}\left(\boldsymbol{W}_{3}\left(\boldsymbol{H}_{1}\odot\boldsymbol{H}_{2}\right)+\boldsymbol{b}_{3}\right)+\boldsymbol{H},$
@@
Meanwhile, due to the linear projection before the RELA, we can treat it as an input gate of linear attention like long-term memory network (LSTM), previous studies (Greff et al., 2017; Van Der Westhuizen and Lasenby, 2018) have demonstrated that the input gate plays a less significant role compared to the forget gate in LSTM. Inspired by this, we remove the linear projection for the value matrix of linear attention. To this end, we only use the weight matrices on query and key, reducing training costs.
3.3. Multi-Layer Perceptron
To capture the complex features, we further leverage a Multi-Layer Perceptron (MLP) (Kang and McAuley, 2018):
(23) $\boldsymbol{H}_{4}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}_{3}\right),$
@@ 3.3. Multi-Layer Perceptron
where $\boldsymbol{W}_{4}\in\mathbb{R}^{D\times 4D},\boldsymbol{W}_{5}\in\mathbb{R}^{4D\times D}$ denote the weight matrices, $\boldsymbol{b}_{4}$, $\boldsymbol{b}_{5}$ are also the biases. $\boldsymbol{H}_{L}$ indicates the final output representation, and $\boldsymbol{H}_{4}$ denotes the output of layer normalization after the GRELA block. To stabilize the training process and enhance gradient flow through the network, we employ a residual connection (He et al., 2016). For generality, the subscript (0) indicates that the final user representation is obtained using a single GRELA layer.
3.4. Prediction Layer and Training

### v2

@@ 3 METHODOLOGY
Figure 2: The overview of our proposed architecture. The Overall Architecture (a) processes an input sequence through an Embedding Layer, which is then fed into a stack of $L$ identical encoder layers. Each encoder layer consists of Layer Normalization, a GRELA Block, another Layer Normalization, and an MLP, finally leading to a Prediction Layer. The central components are the GRELA Block (b), which utilizes RoPE (Rotary Position Encoding, d), and the RELA module (c). Within the RELA module, query ($Q$) and key ($K$) are processed by ELU activation and the RoPE. Attention scores are computed via matrix multiplication of the RoPE-enhanced $Q$ and $K$ (after $K$ is scaled), followed by another scaling and a concatenation step. A Causal Conv1D and SiLU activation processes the value ($V$) projection to capture local context.
@@
In this section, we introduce our proposed framework, RecGRELA. The framework overview is shown in Figure 2. The presentation is structured as follows: initially, we introduce the embedding layer; secondly, we elaborate on the core components of the GRELA Block; subsequently, we present the details of the Multi-Layer Perceptron, and finally, we describe the prediction layer and training strategies.
3.1 Embedding Layer
@@
Following [19], we utilize an embedding layer to map the training sequence $\mathcal{S}_{u}$ into a high-dimensional space. The embedding layer uses a learnable embedding matrix $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $\mathcal{V}$ denotes the item set and $D$ represents the embedding dimension. Given an input interaction sequence $\mathcal{S}_{u}$ of length $N$, it can be projected into the initial embedding matrix $\boldsymbol{H}_{o}\in\mathbb{R}^{N\times D}$ as follows:
$\boldsymbol{H}_{o}=\boldsymbol{E}\mathcal{S}_{u}.$ (9)
@@
Then we apply the Dropout [45] and layer normalization [23] operations to obtain the final embedding $\boldsymbol{H}\in\mathbb{R}^{N\times D}$:
$\boldsymbol{H}=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(\boldsymbol{H}_{o}\right)\right).$ (10)
@@
Attention-based sequential recommendation models [19, 47, 26, 31] typically inject a learnable position encoding into the final embedding, while our method captures position awareness with rotary position encoding by multiplying the rotation matrix with outputs of the kernel function in linear attention. This design eliminates the need to incorporate positional information into the embedding layers.
3.2 GRELA Block
@@
3.2.1 Rotary-Enhanced Linear Attention
@@
Overall Design. Previous work [31] provides $\mathcal{O}(N)$ computational complexity of attention mechanism to alleviate the issue of quadratic complexity in dot-product attention for sequential recommendation. However, it ignores the high cost of learnable position encoding in long-sequence modeling. Meanwhile, some works [50, 34] notice these issues and propose improved RoPE to address them, but unfortunately, they fail to address the problem of dot-product attention with $\mathcal{O}(N^{2})$ complexity. To bridge this gap, we propose a novel Rotary-Enhanced Linear Attention (RELA) mechanism, aiming to simultaneously resolve the issues of quadratic complexity and ineffective position encoding in long-sequence modeling. The motivation stems from the observation that other position encoding methods operate directly on the input sequence or the attention matrix. However, linear attention does not compute the attention matrix, which means other methods can only modify the input sequence. In contrast, RoPE achieves relative position encoding by leveraging absolute position encoding, by applying transformations only to the queries $\boldsymbol{Q}$ and keys $\boldsymbol{K}$. This design makes RoPE suitable for integration with linear attention. In a general situation, following Eq.(7), we revise Eq. (3) as follows:
$\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$ (11)
@@
where $\mathcal{R}\left(\cdot\right)$ denotes injecting the RoPE into queries or keys. Following [20], we set $\phi(x)=\operatorname{ELU}(x)+1$. This kernel function ensures that the denominator remains positive to avoid the risk of division by zero, while the summation in the numerator also contains no negative terms.
@@
Inspired by techniques for stabilizing gradients in dot-product attention [52], we further incorporate scaling factors. Specifically, the keys $\boldsymbol{K}$ and values $\boldsymbol{V}$ are scaled by $\sqrt{N}$ (where $N$ denotes the sequence length). This emulates the scaling mechanism used in conventional dot-product attention and helps regulate the dynamic range of the resulting key-value memory. Furthermore, the normalization term in the denominator, which is critical for linear attention, is constructed using the dot product between the query $\phi(\boldsymbol{Q}_{m})$ and the mean of keys $\phi(\boldsymbol{K}_{n})$. A small constant $\epsilon$ is added to this denominator for numerical stability, preventing division by zero and ensuring well-behaved gradients. This approach enables more fine-grained control over the attention distribution. The final RELA is formulated as follows:
$\boldsymbol{O}_{m}=\frac{\mathcal{R}\left(\phi\left(\boldsymbol{Q}_{m}\right)\right)\left(\left(\mathcal{R}\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)\boldsymbol{V}_{n}/\sqrt{N}\right)\right)}{\phi(\boldsymbol{Q}_{m})\frac{1}{N}\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}+\epsilon}.$ (12)
@@
$\boldsymbol{Q}_{m}=\left[\boldsymbol{Q}_{m}^{(1)},\boldsymbol{Q}_{m}^{(2)}\right]=\left[\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right],$ (13)
@@
$e^{im\theta}=\cos m\theta+i\sin m\theta.$ (14)
@@
$\boldsymbol{Q}_{m}e^{im\theta}$ $=\left(\boldsymbol{Q}_{m}^{(1)}+i\boldsymbol{Q}_{m}^{(2)}\right)(\cos(m\theta)+i\sin(m\theta))$ (15)
@@
$\boldsymbol{Q}_{m}e^{im\theta}=[$ $\boldsymbol{Q}_{m}^{(1)}\cos(m\theta)-\boldsymbol{Q}_{m}^{(2)}\sin(m\theta),$ (16)
$\boldsymbol{Q}_{m}^{(2)}\cos(m\theta)+\boldsymbol{Q}_{m}^{(1)}\sin(m\theta)]$
@@
3.2.2 Local Shortcut and Gated Mechanism
Attention mechanisms excel at long-range modeling, yet local behavior modeling is equally vital for long-term sequential recommendation systems, as user interactions often display strong local preferences (e.g., session-based behaviors) crucial for next-item prediction. However, linear attention variants falter here. Unlike Softmax attention’s sharp focus, linear attention’s smooth distribution tends to average features, obscuring crucial local cues [13]. Moreover, RoPE suffers from long-term decay [46], meaning the attention score decreases as the relative distance increases. Because of linear attention’s smoothing and RoPE’s distance decay, precisely modeling crucial local patterns for long-term sequential recommendation becomes difficult. To address these issues, we introduce a Local Shortcut to mitigate the long-term decay of RELA. Specifically, we use a causal 1D convolution [10] followed by a SiLU activation [41] to capture local preferences in the user interaction sequence:
$\boldsymbol{S}=\operatorname{SiLU}\left(\text{CausalConv1d}\left(\boldsymbol{V}\right)\right),$ (17)
@@ 3.2.2 Local Shortcut and Gated Mechanism
$\boldsymbol{O}_{f}=\boldsymbol{O}+\boldsymbol{S}.$ (18)
@@
The theoretical motivation behind such a design concerns the upper bound on the rank of the attention matrix in linear attention [13]. The rank of the attention matrix is bounded by the length of the sequence $N$ and the feature dimension $d$ for each head:
$\operatorname{rank}(\phi(\boldsymbol{Q})\phi(\boldsymbol{K})^{\top})$ $\leq\min\{\operatorname{rank}(\phi(\boldsymbol{Q})),\operatorname{rank}(\phi(\boldsymbol{K}))\}$ (19)
@@
$\boldsymbol{O}_{f}$ $=\phi(\boldsymbol{Q})\phi(\boldsymbol{K})+\operatorname{CSC}(\boldsymbol{V})$ (20)
@@
To address the challenge of differentiating between temporary local preferences and genuine shifts in long-term user interests, we design a gated mechanism to intelligently modulate RELA’s output. We utilize a SiLU gate [39, 12] and functioning as an output gate, learns to derive weights from the interaction sequence. These weights then adjust the influence of different signals within the user behavior sequence processed by RELA. This allows the model to discern whether a new burst of user behaviors represents a strong but potentially transient local preference or signifies a true evolution in the user’s underlying taste. The overall structure of the GRELA Block is shown as follows:
$\boldsymbol{H}_{g}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}\right),$ (21)
@@
where $\boldsymbol{H}_{g}$ is the input of the GRELA block, i.e., when $L=1$, the input $\boldsymbol{H}$ is from the embedding layer, $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}\in\mathbb{R}^{D\times D}$ are the weight matrices of the linear projection, $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$ are the biases. Then a Hadamard product is performed on $\boldsymbol{H}_{1}$ and $\boldsymbol{H}_{2}$ followed by a linear projection and Droppath [22] operation:
$\boldsymbol{H}_{3}=\operatorname{DropPath}\left(\boldsymbol{W}_{3}\left(\boldsymbol{H}_{1}\odot\boldsymbol{H}_{2}\right)+\boldsymbol{b}_{3}\right)+\boldsymbol{H},$ (22)
@@
Meanwhile, due to the linear projection before the RELA, we can treat it as an input gate of linear attention like long-term memory network (LSTM), previous studies [11, 51] have demonstrated that the input gate plays a less significant role compared to the forget gate in LSTM. Inspired by this, we remove the linear projection for the value matrix of linear attention. To this end, we only use the weight matrices on query and key, reducing training costs.
3.3 Multi-Layer Perceptron
To capture the complex features, we further leverage a Multi-Layer Perceptron (MLP) [19]:
$\boldsymbol{H}_{4}$ $=\operatorname{LayerNorm}\left(\boldsymbol{H}_{3}\right),$ (23)
@@ 3.3 Multi-Layer Perceptron
where $\boldsymbol{W}_{4}\in\mathbb{R}^{D\times 4D},\boldsymbol{W}_{5}\in\mathbb{R}^{4D\times D}$ denote the weight matrices, $\boldsymbol{b}_{4}$, $\boldsymbol{b}_{5}$ are also the biases. $\boldsymbol{H}_{L}$ indicates the final output representation, and $\boldsymbol{H}_{4}$ denotes the output of layer normalization after the GRELA block. To stabilize the training process and enhance gradient flow through the network, we employ a residual connection [16]. For generality, the subscript (0) indicates that the final user representation is obtained using a single GRELA layer.
3.4 Prediction Layer and Training